OpenStudy (anonymous):

do you combine 10b^2 and 3a^2 or just leave it alone? its a part of polynomial question

jimthompson5910 (jim_thompson5910):

they are not like terms or common terms, so you cannot combine them if you had 10b^2 and 3b^2, then you could combine them to get 10b^2+3b^2 = (10+3)b^2 = 13b^2 however, the second term has 'a^2' in it which means the two terms aren't like terms

OpenStudy (bibby):

you can make the idea abstract and think about adding apples and oranges. a^2 + a^2 = 2a^2 a^2 + b^2 = a^2 + b^2 (you can't add apples and oranges)